OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -3 f(1) = 6 f(n) = -2•f(n -1) - f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -3, 6, -9, -12, … -3, 20, -95, 480, … -3, 6, -9, 12, … -3, -20, -95, -480, …

OpenStudy (anonymous):

i will medal the most helpful person

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

i need help

hero (hero):

I see

hero (hero):

Have you attempted to find f(2)?

OpenStudy (anonymous):

f(2)=15?

OpenStudy (anonymous):

@Hero f(2)=15?

hero (hero):

f(2) = -2•f(2 -1) - f(2 - 2); for n > 1 f(2) = -2f(1) - f(0) Hint: f(1) and f(0) are given.

hero (hero):

2nd Hint: f(2) = -2(6) - (-3) = -12 + 3 = -9

OpenStudy (anonymous):

so f(2)=-9?

hero (hero):

Correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that narrows it down to 2 answers

hero (hero):

That narrows down the answer choices.

hero (hero):

Beat me to it.

OpenStudy (anonymous):

the first one and the third one lol

hero (hero):

Actually, it narrows it down to one choice if you notice the pattern.

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

the third one

hero (hero):

Correct

OpenStudy (anonymous):

thanks